OpenStudy (scarlettfarra2000):

Working with Polynomials Multiply and simplify. (a) 〖3x〗^2 y^8∙7x^5 y^(-4) (b) (x+3)(x-5) (c) (x+4)^2 Simplify and leave answers in STANDARD FORM. (a) (9x^2+8x-5)-(7x^2+2x-2) (b) (9x^2+8x-5)-(7x^2+2x-2)

OpenStudy (qwertty123):

Do you know how to start these?

OpenStudy (scarlettfarra2000):

No

OpenStudy (qwertty123):

Do you have to do all of these?

OpenStudy (scarlettfarra2000):

Yep

OpenStudy (qwertty123):

B.) (x+3)(x-5) =(x+3)(x+−5) =(x)(x)+(x)(−5)+(3)(x)+(3)(−5) =x2−5x+3x−15

OpenStudy (qwertty123):

@imqwerty Is great at math he can help you with the rest! I have to go very soon.

OpenStudy (scarlettfarra2000):

thanks you for the help

OpenStudy (scarlettfarra2000):

@imqwerty are you going to help? I really need it

OpenStudy (qwertty123):

lol It shows he is here but I don't know?

OpenStudy (scarlettfarra2000):

neither do I

OpenStudy (scarlettfarra2000):

@satellite73 would you please help if you know how to do this